What is the heat capacity of gold if a 250 g sample needs 133 cal of energy to increase its temperature from 23.0 degrees Celsius to 40.1 degrees Celsius
1
Expert's answer
2019-11-21T06:26:00-0500

Q=c*m*(t2-t1); 

c=Q/[m*(t2-t1)]=133 cal/[0.25 kg*(40.1-23.0)C]=31.1 cal/(kg*C)